### 20 Dec 2023 v1

# A faster FPRAS for #NFA

Abstract.
Given a non-deterministic finite automaton (NFA) $\mathcal{A}$ with $m$ states, and a natural number $n\in\mathbb{N}$ (presented in unary), the #NFA problem asks to determine the size of the set $\mathcal{L}(\mathcal{A}_{n})$ of words of length $n$ accepted by $\mathcal{A}$. While the corresponding decision problem of checking the emptiness of $\mathcal{L}(\mathcal{A}_{n})$ is solvable in polynomial time, the #NFA problem is known to be #P-hard. Recently, the long-standing open question — whether there is an FPRAS (fully polynomial time randomized approximation scheme) for #NFA — was resolved in (Arenas et al., 2019). The FPRAS due to (Arenas et al., 2019) relies on the interreducibility of counting and sampling, and computes, for each pair of state $q$ and natural number $i\leq n$, a set of $O(\frac{m^{7}n^{7}}{\epsilon^{7}})$ many uniformly chosen samples from the set of words of length $i$ that have a run ending at $q$. This informative measure — the number of samples maintained per state and length — also affects the overall time complexity with a quadratic dependence. Given the prohibitively high time complexity, in terms of each of the input parameters, of the FPRAS due to (Arenas et al., 2019), and considering the widespread application of approximate counting (and sampling) in various tasks in Computer Science, a natural question arises: Is there a faster FPRAS for #NFA that can pave the way for the practical implementation of approximate #NFA tools? In this work, we demonstrate that significant improvements in time complexity are achievable. Specifically, we have reduced the number of samples required for each state to be independent of $m$, with significantly less dependence on $n$ and $\epsilon$, maintaining only $\widetilde{O}(\frac{n^{4}}{\epsilon^{2}})$ samples per state. We believe our insights may lead to further reductions in time complexity and thus open up a promising avenue for future work towards the practical implementation of tools for approximating #NFA.
The authors decided to forgo the old convention of alphabetical ordering of authors in favor of a randomized ordering, denoted by r⃝. The publicly verifiable record of the randomization is available at https://www.aeaweb.org/journals/policies/random-author-order/search
copyright: noneconference: ; ;
1. Introduction
In this paper, we focus on the following computational problem:
  #NFA:
: Given an NFA $\mathcal{A}=(Q,I,\Delta,F)$ over the binary alphabet $\Sigma=\{0,1\}$ with $m$ states, and a number $n$ in unary, determine $|\mathcal{L}(\mathcal{A}_{n})|$, where $\mathcal{L}(\mathcal{A}_{n})$ is the set of strings of length $n$ that are accepted by $\mathcal{A}$.
The counting problem #NFA and the closely related problem of uniform sampling words of an NFA are fundamental to database applications, and arise when counting and uniformly generating solutions to queries. Beyond database systems, the aforementioned problem finds applications in program verification (Legay et al., 2010), testing of software (Sutton et al., 2007) and distributed systems (Ozkan et al., 2019), probabilistic programming, estimation of information leakage in software (Bang et al., 2016; Gao et al., 2019; Saha et al., 2023), and even music generation (Donzé et al., 2014).
How do we count the number of elements of $\mathcal{L}(\mathcal{A}_{n})$? When $\mathcal{A}$ is a DFA with $m$ states, this question can be solved using a simple dynamic programming algorithm that inductively computes, for every $i\leq n$ and for every state $q$ of $\mathcal{A}$, the number of words of length $i$ that end in state $q$: $\mathsf{count}(q,i)=\sum_{b\in\Sigma}\sum_{p:q=\Delta(p,b)}\mathsf{count}(p,i-1)$ taking overall $O(mn)$ time (assuming $|\Sigma|=O(1)$). This simple algorithm however fails to work when $\mathcal{A}$ is a non-deterministic automaton as the same word maybe accepted using many different runs. Determinizing $\mathcal{A}$ first and then using the same algorithm as for DFAs would give an overall running time of $O(n\cdot 2^{m})$ which is exponential in the size of the description of the input. It turns out that #NFA is, in fact, provably computationally intractable — it is known to be a #P-hard problem (Álvarez and Jenner, 1993).
Given the intractability, and given the widespread applications of this problem, the next natural question then arises — can we approximate #NFA using a polynomial time algorithm? One of the promising paradigms for deriving efficient algorithms for approximating the answer to an intractable problem is the use of randomization to obtain a fully polynomial time randomized approximation scheme (FPRAS). Formally, an FPRAS for a real-valued function $f:\mathcal{D}\to\mathbb{N}$ over some domain $\mathcal{D}$ is a randomized algorithm $A:\mathcal{D}\times(0,1)\times(0,1)\to\mathbb{N}$ that for every input $x\in\mathcal{D}$ and parameters $\epsilon\in(0,1)$ and $\delta\in(0,1)$, runs in time $\textsf{poly}(|x|,\frac{1}{\epsilon},\log(\frac{1}{\delta}))$, and satisfies: ${\mathsf{Pr}}\!\left[\frac{1}{(1+\epsilon)}\cdot f(x)\leq|A(x,\epsilon,\delta)|\leq(1+\epsilon)\cdot f(x)\right]\geq 1-\delta$.
Until very recently, the only approximation algorithms known for #NFA had running times that were quasi-polynomial in the size of the inputs (Gore et al., 1997; Kannan et al., 1995). In a recent breakthrough by Arenas, Croquevielle, Jayram and Riveros (Arenas et al., 2019, 2021), #NFA was finally shown to admit an FPRAS. Even though the problem was shown to admit an FPRAS, the time complexity of the proposed FPRAS is prohibitively large — it takes time $\widetilde{O}\left(m^{17}n^{17}\cdot\frac{1}{\epsilon^{14}}\cdot\log(1/\delta)\right)$, where the notation $\widetilde{O}$ hides the polylogarithmic multiplicative factors in $m$ and $n$. To be applicable in a practical setting like estimating information flow in programs, a polynomial time approximation scheme whose time complexity varies as $O(|P|^{16})$, where $|P|$ is roughly the lines of code, is unlikely to scale even for moderately sized programs. Towards this, in this work, we pursue the goal of desigining a faster algorithm, and propose an FPRAS which runs in time $\widetilde{O}(m^{4}n^{11}\cdot\frac{1}{\epsilon^{4}}\cdot\log^{2}(1/\delta))$. The reliance of both FPRASes on the inter-reducibility of sampling and counting necessitates a more informative measure, such as the number of samples maintained for each state. The scheme proposed in  (Arenas et al., 2021) maintains $O(\frac{m^{7}n^{7}}{\epsilon^{7}})$ samples per state while the scheme proposed in this work maintains only $O(\frac{n^{4}}{\epsilon^{2}})$ samples, which is independent of $m$ and with a significant reduction in it’s dependence on $n$ and $\epsilon$.
1.1. Overview of our FPRAS
At a high-level, our approach exploits the tight connection between approximate counting and the related problem of (almost) uniform generation. In the context of an NFA $\mathcal{A}$ and a length $n$, the almost uniform generation problem asks for a randomized algorithm (generator) $G$, such that, on each invocation, the probability with which $G$ returns any given string $x\in\mathcal{L}(\mathcal{A}_{n})$ lies in the range $[\frac{1}{|\mathcal{L}(\mathcal{A}_{n})|}-\epsilon,\frac{1}{|\mathcal{L}(\mathcal{A}_{n})|}+\epsilon]$; here $\epsilon\in(0,1)$ is a parameter. Such a random generator is a polynomial time generator if its running time is polynomial in $|\mathcal{A}|$ and $n$. The existence of a polynomial time almost uniform generator for self-reducible (Jerrum et al., 1986) problems implies the existence of an FPRAS, and vice versa.
Input : NFA $\mathcal{A}$ with $m$ states $Q$, $n\in\mathbb{N}$ (in unary)
1 Unroll the automaton $\mathcal{A}$ into an acyclic graph $\mathcal{A}_{\textsf{unroll}}$ by making $n+1$ copies $\{q^{\ell}\}_{\ell\in[0,n],q\in Q}$ of the states $Q$ and adding transitions between immediate layers foreach $\ell\in\{0,\ldots,n\},q\in Q$ do
2    Compute the estimate $\mathsf{N}({q}^{\ell})$ using prior estimates and samples $\{\mathsf{N}({q}^{i}),\mathsf{S}({q}^{i})\}_{i<\ell,p\in Q}$ Compute a uniform set of samples $\mathsf{S}({q}^{\ell})$ using $\mathsf{N}({q}^{\ell})$ and the prior estimates and samples $\{\mathsf{N}({q}^{i}),\mathsf{S}({q}^{i})\}_{i<\ell,p\in Q}$
return $\mathsf{N}({F}^{\ell})$
Algorithm 1 Algorithm Template: FPRAS for #NFA
This tight connection between counting and sampling was also the high-level idea behind the work of (Arenas et al., 2021). Consequently, our algorithm fits the template procedure of Algorithm 1, also outlined in (Arenas et al., 2021). For technical convenience, the algorithm unrolls the automaton $\mathcal{A}$ into an acyclic graph with $n$ levels; each level $\ell$ containing the $\ell^{\text{th}}$ copy ${q}^{\ell}$ of each state $q$ of $\mathcal{A}$. A key component in the algorithm is to incrementally compute an estimate $\mathsf{N}({q}^{\ell})$ of the size of the set $\mathcal{L}({q}^{\ell})$ of words of length $\ell$ that have a run ending in state $q$.
Next, in order to inductively estimate $\mathsf{N}({q}^{\ell})$, we in turn rely on the estimates of sets $\mathcal{L}({p_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}}}}^{\ell-1})$ (${\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}}\in\Sigma=\{0,1\}$); here $p_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}}}$ is a ${\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}}$-predecessor state of $q$, i.e., $(p_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}}},{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}},q)\in\Delta$ is a transition of $\mathcal{A}$. These estimates can be put to use effectively because:
$\mathcal{L}({q}^{\ell})=\big{(}\bigcup_{(p_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{0}}},{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{0}},q)\in\Delta}\mathcal{L}({p_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{0}}}}^{\ell-1})\big{)}\cdot\{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{0}}\}\cup\big{(}\bigcup_{(p_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{1}}},{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{1}},q)\in\Delta}\mathcal{L}({p_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{1}}}}^{\ell-1})\big{)}\cdot\{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{1}}\}.$
We remark that $\mathcal{L}({p}^{\ell-1})$ may not be disjoint for different predecessor states $p$. Hence, we cannot estimate $\bigcup_{(p_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}}},{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}},q)\in\Delta}\mathcal{L}({p_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}}}}^{\ell-1})$ by merely summing the individual estimates $\{\mathsf{N}({p_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}}}}^{\ell-1})\,|\,(p_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}}},{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}},q)\in\Delta\}$. To address this, we maintain a set, $\mathsf{S}({p}^{\ell-1})$, for each $p$, which contains samples of strings from $\mathcal{L}({p}^{\ell-1})$. One possible idea is to construct the set $\mathsf{S}({p}^{\ell-1})$ using strings $w$ that have been sampled uniformly at random from $\mathcal{L}({p}^{\ell-1})$ but such a process introduces dependence among the sets $\mathsf{S}({p}^{\ell-1})$ for different $p$. The crucial observation in (Arenas et al., 2021) is that for any $k$, the following is true: for any word $w$ of length $k$, all words from $\mathcal{L}({q}^{\ell})$ with the suffix $w$ can be expressed as $\big{(}\bigcup\limits_{p\in P}\mathcal{L}({p}^{\ell-k})\big{)}\cdot\{w\}$ for some set $P$ of states, where $P$ depends only on $w,q,$ and $k$. This property is referred to as the self-reducible union property. This insight allows us to utilize the self-reducibility property for sampling. We can partition $\mathcal{L}(p^{\ell-1})$ into sets of strings where the last character is either ${\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{0}}$ or ${\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{1}}$, and compute estimates for each partition since each partition can be represented as a union of sets for which we (inductively) have access to both samples and estimates. We then sample the final character in proportion to the size of these estimates. In a similar recursive manner, we can sample strings, character by character. That is, we can sample strings by progressively extending suffixes of $w$ backwards! Of course, we must also account for the errors that propagate in the estimates of $\mathsf{N}({q}^{\ell})$.
Our algorithm also relies on the self-reducible union property, and therefore, the basic structure of our algorithm is similar to that of (Arenas et al., 2021) on high-level. There are, however, significant and crucial technical differences that allow us to obtain an FPRAS with significantly lower time complexity. We highlight these differences below. As a first step, we highlight the two key invariants underpinning the algorithm and ensuing analysis in (Arenas et al., 2021):
  1. (1)
    The following condition (Inv-ACJR) holds for every level $\ell$ (here, $\kappa=\frac{nm}{\epsilon}$), with high probability.
    (Inv-ACJR) $\begin{array}[]{l}\xi(\ell)\triangleq\forall q{\in}Q,\forall P{\subseteq}Q,\big{|}\frac{|\mathsf{S}({q}^{\ell})\setminus\bigcup\limits_{p\in P}\mathcal{L}(p^{\ell})|}{|\mathsf{S}({q}^{\ell})|}-\frac{|\mathcal{L}(q^{\ell})\setminus\bigcup\limits_{p\in P}\mathcal{L}(p^{\ell})|}{|\mathcal{L}(q^{\ell})|}\big{|}\leq\frac{1}{\kappa^{3}}\end{array}$
  2. (2)
    $\mathsf{S}({q}^{\ell})$ is close in $(L_{\infty}$ distance)111 In (Arenas et al., 2021), the invariant is stated in a different, but equivalent, formulation: conditioned on $\xi(\ell)$, the set $\mathsf{S}({q}^{\ell})$ follows the distribution of a multi-set obtained by independently sampling every element of $\mathcal{L}(q^{\ell})$ uniformly at random, with replacement. to the distribution of multi-set obtained by independently sampling (with replacement) every element of $\mathcal{L}(q^{\ell})$ uniformly at random.
In contrast, the FPRAS we propose maintains the following two weaker invariants:
  1. (1)
    For every state $q$ and level $\ell$, the event that $\mathsf{N}({q}^{\ell})\in(1\pm\frac{\epsilon}{2n^{2}})|\mathcal{L}(q^{\ell})|$, which we denote as $\texttt{AccurateN}_{q,\ell}$, happens with high probability.
  2. (2)
    The distribution of $\mathsf{S}({q}^{\ell})$ is close, in total variational distance, to the distribution of a multi-set constructed by sampling (with replacement) every element of $\mathcal{L}(q^{\ell})$ uniformly at random.
Two remarks are in order. First, $\xi(\ell)$ implies $\texttt{AccurateN}_{q,\ell}$, and closeness in $L_{\infty}$ is a stringent condition than closeness in total variation distance. Second, the technical arguments of  (Arenas et al., 2021) crucially rely on their invariants and do not work with the aforementioned weaker arguments. Accordingly, this necessitates a significantly different analysis that depends on coupling of the distributions. It’s also noteworthy that, due to our reliance on the weaker condition, our method for estimating $\mathsf{N}({q}^{\ell})$ differs from that in (Arenas et al., 2021). In fact, our approach is based on an adaptation of the classical Monte Carlo-based computation of the union of sets (Karp and Luby, 1985).
For a given state $q$ and subset $P$, achieving a $\frac{1}{\kappa^{3}}$-additive approximation for $\frac{|\mathcal{L}(q^{\ell})\setminus\cup_{p\in P}\mathcal{L}(p^{\ell})|}{|\mathcal{L}(q^{\ell})|}$ with a probability of $1-\eta$ requires, according to standard concentration bounds, that the size of $\mathsf{S}({q}^{\ell})$ be $O(\kappa^{6}\log\eta^{-1})$. To ensure this additive approximation for all states $q$ and subsets $P$, one must consider union bounds over exponentially many events. Specifically, for the analysis to be successful, one must have $\eta^{-1}\in O(2^{mn})$. Consequently, the FPRAS of (Arenas et al., 2021) maintains $O(\kappa^{7})$ samples for each state $q$. In contrast, for $\texttt{AccurateN}_{q,\ell}$ to hold with a probability of $1-\eta$, our FPRAS only maintains $O\left(\frac{n^{4}}{\epsilon^{2}}\right)\log\eta^{-1}$ samples in $\mathsf{S}({q}^{\ell})$. Further, a union bound over only $O(mn)$ events is required, thus setting $\eta^{-1}=O(mn)$ is adequate.
The time complexity of an algorithm that follows the template outlined in Algorithm 1 (including both the FPRAS of (Arenas et al., 2021) as well as the FPRAS developed in this paper), can be expressed as $O(n\cdot(m\cdot n)\cdot(m^{2}n)\cdot|\mathsf{S}({q}^{\ell})|^{2})$ where the first $n$ from the length of every string, the factor $(m\cdot n)$ corresponds to the upper bound on the number of states in the unrolled automaton, and the factor $(m^{2}n)$ corresponds to the time taken for membership check. Accordingly, it is instructive to compare the bound of $O(\frac{m^{7}n^{7}}{\epsilon^{7}})$ for $|S(q^{\ell})|$ in  (Arenas et al., 2021) to $O(\frac{n^{4}}{\epsilon^{2}})$ in our approach. Remarkably, the number of samples for every state in our approach is independent of $m$, the size of the automaton.
2. Preliminaries
We consider the binary alphabet $\Sigma=\{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{0}},{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{1}}\}$. All results in this paper are extendable to alphabets of arbitrary but fixed constant size. A string $w$ over $\Sigma$ is either the empty string $\lambda$ (of length 00) or a non-empty finite sequence $w_{1}w_{2}\ldots w_{k}$ (of length $|w|=k>0$) where each $w_{i}\in\Sigma$. The set of all strings over $\Sigma$ is denoted by $\Sigma^{*}$. A non-deterministic finite automaton is a tuple $\mathcal{A}=(Q,I,\Delta,F)$ where $Q$ is a finite set of states, $I\in Q$ is the initial state, $\Delta\subseteq Q\times\Sigma\times Q$ is the transition relation, and $F$ is the set of accepting states. A run of $w$ on $\mathcal{A}$ is a sequence $\rho=q_{0},q_{1}\ldots,q_{k}$ such that $k=|w|$, $q_{0}=I$ and for every $i<k$, $(q_{i},w_{i+1},q_{i+1})\in\Delta$; $\rho$ is accepting if $q_{k}\in F$. The word $w$ is accepted by $\mathcal{A}$ if there is an accepting run of $w$ on $\mathcal{A}$. The set of strings accepted by $\mathcal{A}$ is $\mathcal{L}(\mathcal{A})$. For a natural number $n\in\mathbb{N}$, the $n^{\text{th}}$ slice of $\mathcal{A}$’s language, $\mathcal{L}(\mathcal{A}_{n})$, is the set of strings of length $n$ in $\mathcal{L}(\mathcal{A})$. For simplicity, this paper assumes a singleton set of accepting states, i.e., $F=\{q_{F}\}$ for some $q_{F}\in Q$. We introduce additional notation for our algorithms. For $q\in Q$ and $b\in\Sigma$, the $b$-predecessors of $q$ are:
$\textsf{Pred}(q,b)=\{p\,|\,(p,b,q)\in\Delta\}.$
For each $q\in Q$ and $1\leq\ell\leq n$, $\mathcal{L}(q^{\ell})$ denotes the set of strings $w$ of length $\ell$ with a run on $\mathcal{A}$ ending in $q$. The rest of the paper focuses on the automaton $\mathcal{A}=(Q,I,\Delta,F)$, and a number $n$ in unary, using $m$ to denote $|Q|$.
3. A Faster FPRAS
Given automaton $\mathcal{A}$ and a number $n$ in unary, we construct the unrolled automaton $\mathcal{A}_{\textsf{unroll}}$. We use the notation ${q}^{\ell}$ to denote the $\ell^{\text{th}}$ copy of state $q\in Q$. We use $\mathcal{L}({q}^{\ell})$ to denote the the set of all distinct strings $w$ of length $\ell$ for which there is a path (labeled by $w$) from the starting state to $q$ (i.e., those strings that go to state ${q}^{\ell}$ in $\mathcal{A}_{\textsf{unroll}}$). We will use $\mathsf{N}({q}^{\ell})$ to denote an estimate of $|\mathcal{L}({q}^{\ell})|$.
In Section 3.1, we formally present our FPRAS for #NFA, together with the statement of its correctness and running time (Theorem 3). Our FPRAS Algorithm 4 calls two subroutines Algorithm 3 and Algorithm 2, each of which are presented next, together with their informal descriptions. In later sections, we present the formal proofs of their correctness.
3.1. Description of the Algorithms
Algorithm 4 presents the main procedure for approximating the size of $\mathcal{L}(\mathcal{A})_{n}$, where $\mathcal{A}$ is the NFA and $n$ is the parameter that is given as input. The algorithm works in a dynamic programming fashion and maintains, for each state $q$ and each length $1\leq\ell\leq n$, a sampled subset $\mathsf{S}({q}^{\ell})$ of $\mathcal{L}({q}^{\ell}))$. As a key step, it uses the sampled sets corresponding to $\textsf{Pred}(q,{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{0}})$ and $\textsf{Pred}(q,{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{1}})$ to construct the desired sampled subset of $\mathcal{L}({q}^{\ell}))$. We first describe the two subroutines the $\texttt{AppUnion}{()}$ and $\mathsf{sample}()$. We start with describing the subroutines first.
1 function AppUnion ${}_{(\epsilon,\delta)}$($\epsilon_{\textsf{sz}},(O_{1},S_{1},\textsf{sz}_{1}),\ldots,(O_{k},S_{k},\textsf{sz}_{k})$)
2     $m\leftarrow\lceil\frac{\sum_{i}\textsf{sz}_{i}}{\max_{i}\textsf{sz}_{i}}\rceil$ $t\leftarrow\frac{12\cdot(1+\epsilon_{\textsf{sz}})^{2}\cdot m}{\varepsilon^{2}}\cdot\log(\frac{4}{\delta})$ $Y\leftarrow 0$ repeat $t$ times:
3          With prob. $p_{i}=\frac{\textsf{sz}_{i}}{\sum_{i}\textsf{sz}_{i}}$, sample $i\in\{1,\ldots,k\}$ if $S_{i}\neq\varnothing$ then
4               $\sigma\leftarrow S_{\ell}\!\boldsymbol{\cdot}\!\mathsf{dequeue}()$
5         else
6               break
7         if $\not\exists j<i,$ such that $O_{j}\!\boldsymbol{\cdot}\!\mathsf{contains}(\sigma)$ then  $Y\leftarrow Y+1$
8    return $\frac{Y}{t}\cdot{\sum_{i}\textsf{sz}_{i}}$
Algorithm 2 Approximating Union of Sets
1 function $\ell$, ${P}^{\ell}$, $w$, $\varphi$,$\beta$, $\eta$)
2     $\eta^{\prime}\leftarrow\eta/4n$ $\beta^{\prime}\leftarrow(1+\beta)^{\ell-1}-1$ if $\ell=0$ then
3          if $\varphi>1$ then return $\bot$ w.p.  ($\varphi$) : return $w$ ||  w.p.  ($1-\varphi$) : return $\bot$
4    else
5          foreach ${\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}}\in\{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{0}},{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{1}}\}$ do
6               ${P}^{\ell}_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}}}\leftarrow\bigcup_{p\in{P}^{\ell}}\textsf{Pred}(p,{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}})$ // From the above set, remove all states that are not reachable from the initial state.
[…]
8
9         $\mathsf{pr}_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{0}}}\leftarrow\frac{\displaystyle\textsf{sz}_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{0}}}}{\displaystyle\textsf{sz}_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{0}}}+\textsf{sz}_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{1}}}}$ w.p.  ($\mathsf{pr}_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{0}}}$) : ${\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}}\leftarrow{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{0}}$ ||  w.p.  ($1-\mathsf{pr}_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{0}}}$) : ${\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}}\leftarrow{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{1}}$ ${P}^{\ell-1}\leftarrow{P}^{\ell}_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}}}$ $w\leftarrow{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}}\cdot w$
10    return sample($\ell-1$, ${P}^{\ell-1}$, $w$, $\varphi/\mathsf{pr}_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}}},\beta,\eta$)
Algorithm 3 Sampling subroutine for $\mathcal{A}_{\textsf{unroll}}$, length $n$
The AppUnion subroutine. AppUnion ( Algorithm 2) takes as input $\epsilon,\delta$ and $\epsilon_{\textsf{sz}}$ and assumes access to sets $T_{1},\dots,T_{k}$. For each set $T_{i}$ the access is given in the form of membership oracle $O_{i}$, a subset $S_{i}$ (obtained by sampling $T_{i}$ with replacements), and an estimate $\textsf{sz}_{i}$ of the size of $T_{i}$. Using these, Algorithm 2 outputs an $(\epsilon,\delta)$ estimate of the size of the union $\cup_{i=1}^{k}T_{i}$.
The algorithm presented herein represents a modification of classical Monte Carlo-based scheme due to Karp and Luby (Karp and Luby, 1985). The proof methodology also shares similarities with  (Karp and Luby, 1985). For an arbitrary $\sigma\in\cup_{i=1}^{k}T_{i}$, let $L(\sigma)$ denote the smallest index $i$ such that $\sigma\in T_{i}$. And let set $\mathcal{U}_{1}$ be the set of all the $(\sigma,L(\sigma))$ pairs and let $\mathcal{U}_{2}$ be the set of all $(\sigma,i)$ pairs such that $\sigma\in T_{i}$. Consequently, the cardinality of the set $\mathcal{U}_{1}$ is $|\cup_{i=1}^{k}T_{i}|$, and that of $\mathcal{U}_{2}$ is $\sum_{i=1}^{k}|T_{i}|$. By drawing a sufficient number of samples from $\mathcal{U}_{2}$ and assessing the fraction belonging to $\mathcal{U}_{1}$, one can estimate $|\mathcal{U}_{1}|/|\mathcal{U}_{2}|$, which when multiplied by $|\mathcal{U}_{2}|$ provides a reliable estimate for the size of $\mathcal{U}_{1}$.
In the algorithm we take $t$ samples from $\mathcal{U}_{2}$. This involves initially sampling a set $T_{i}$ proportional to the size of the set (in line 2 ) and subsequently uniformly drawing a sample from $T_{i}$ (in line 2). As the exact sizes of the sets $T_{i}$ are unknown , estimates $\textsf{sz}_{i}$ are utilized for proportional sampling (in line 2), introducing an additional factor of error in the final estimate of $|\mathcal{U}_{1}|/|\mathcal{U}_{2}|$.This error can be mitigated by adjusting the value of $t$ accordingly.
Upon sampling an $i$ in Line 2, an element is chosen from the sample of the sets $S_{i}$. This mimics a random sample from $T_{i}$, given that $S_{i}$ already contains uniform samples from $T_{i}$. For each sampled index $i$ , a new element from $S_{i}$ is used and removed (Line 2). Initially, $S_{i}$ are sets are guaranteed to have a size of at least thresh, surpassing the expected number of samples required from $T_{i}$ during the $t$ iterations of the loop. If the algorithm ever needs more samples than $|\mathcal{U}_{i}|$ during the $t$ iterations, it counts as an error, but the probability of this occurrence is demonstrated to be very low.
Finally, line 2 verifies whether the sample from $\mathcal{U}_{2}$ belongs to $\mathcal{U}_{1}$. It uses the oracles to the sets to check this. The variable $Y$ tallies the number of samples from $\mathcal{U}_{1}$, and after $t$ iterations, $Y/t$ provides a estimation for $|\mathcal{U}_{1}|/|\mathcal{U}_{2}|$. The final output is the product of this value with $(\sum_{i}\textsf{sz}_{i})$, the estimate for $|\mathcal{U}_{2}|$.
Theorem 1 characterizes the formal guarantee of Algorithm 2.
Theorem 1.
Let $\epsilon,\delta>0$, and let $\Omega$ be some set. Let $T_{1},T_{2},\ldots,T_{k}\subseteq\Omega$ be sets with membership oracles $O_{1},\ldots,O_{k}$ respectively. Let $\epsilon_{\textsf{sz}}\geq 0$ and let $\textsf{sz}_{1},\ldots,\textsf{sz}_{k}\in\mathbb{N}$ be such that for every $i\leq k$, we have $\frac{\left|T_{i}\right|}{(1+\epsilon_{\textsf{sz}})}\leq\textsf{sz}_{i}\leq(1+\epsilon_{\textsf{sz}})\left|T_{i}\right|$. For each $i\leq k$, let $S_{i}$ be a sequence of samples (with possible repetitions) obtained by repeatedly sampling elements from $T_{i}$ uniformly with replacement such that $|S_{i}|\geq\texttt{thresh}$, where $\texttt{thresh}=24\cdot\frac{(1+\epsilon_{\textsf{sz}})^{2}}{\epsilon^{2}}\cdot\log(\frac{4k}{\delta})$ Then, on input $\epsilon_{\textsf{sz}},(O_{1},S_{1},\textsf{sz}_{1}),\ldots,(O_{k},S_{k},\textsf{sz}_{k})$, the algorithm AppUnion with parameters $\epsilon,\delta$ (Algorithm 2) outputs sz that satisfies:
${\mathsf{Pr}}\!\left[\frac{|\cup_{i=1}^{k}T_{i}|}{(1+\epsilon)(1+\epsilon_{\textsf{sz}})}\leq\left|\textsf{sz}\right|<(1+\epsilon)(1+\epsilon_{\textsf{sz}})|\cup_{i=1}^{k}T_{i}|\right]>1-\delta.$
The worst case number of calls to the oracles $O_{i}$ made by the algorithm is $O\left(k\cdot(1+\epsilon_{\textsf{sz}})^{2}\cdot\frac{1}{\epsilon^{2}}\cdot\log\big{(}\frac{k}{\delta}\big{)}\right)$ and the rest of the computation takes a similar amount of time.
[…]
We next tread towards formally characterizing the guarantee of Algorithm 3. To this end, we will introduce few notations:
  1. (1)
    For each $1\leq i\leq\ell$ and each $q\in Q$, let $\texttt{AccurateN}_{i,q}$ be the event that $\frac{|\mathcal{L}({q}^{i})|}{(1+\beta)^{i}}\leq\mathsf{N}({q}^{i})\leq(1+\beta)^{i}|\mathcal{L}({q}^{i})|$. Also, let $\texttt{AccurateN}_{i}=\bigcap\limits_{q\in Q}\texttt{AccurateN}_{i,q}$ and $\texttt{AccurateN}_{\leq i}=\bigcap\limits_{1\leq j\leq i}\texttt{AccurateN}_{j}$.
  2. (2)
    For a state ${q}^{\ell}$, we will use $\mathsf{U}({q}^{\ell})$ to denote the random variable denoting the sequence of samples (with possible repetitions) constructed by repeatedly sampling $|\mathsf{S}({q}^{\ell})|$ number of elements uniformly from $\mathcal{L}({q}^{\ell})$. Our algorithms compute sequences $\mathsf{S}({q}^{\ell})$ of elements (strings) sampled from $\mathcal{L}({q}^{\ell})$. For arguing about correctness of our algorithms, we will in turn argue that the distribution of $\mathsf{S}({q}^{\ell})$ would be “similar" to that of $\mathsf{U}({q}^{\ell})$. Towards this, we will often invoke the the probability space obtained by replacing the sets $\mathsf{S}({q}^{j})$ in Algorithm 3 and Algorithm 4 by $\mathsf{U}({q}^{j})$ for all $q\in Q$ and for all $1\leq j\leq i$. Formally, for any event $\mathcal{E},\mathcal{F}$ (in the context of Algorithm 3 and Algorithm 4) we will denote by ${\mathsf{Pr}}\!\left[\mathcal{E};{{\mathsf{U}}_{\leq i}}\right]$ (resp. ${\mathsf{Pr}}\!\left[\mathcal{E}\;|\;\mathcal{F};{{\mathsf{U}}_{\leq i}}\right]$) the probability of the event $\mathcal{E}$ (resp. probability of the event $\mathcal{E}$, given event $\mathcal{F}$) when the $\mathsf{S}({q}^{j})$ in Algorithm 3 and Algorithm 4 is replaced by $\mathsf{U}({q}^{j})$ for all $q\in Q$ and for all $1\leq j\leq i$.
  3. (3)
    Consider the random trial corresponding to the call to the procedure $\textnormal{{sample}}(\ell,\{{q}^{\ell}\},\lambda,\gamma_{0},\beta,\eta)$ for $\ell>0$ and let $\mathtt{ret}_{\textnormal{{sample}}}$ denote the random variable representing the return value thus obtained. Observe that, in each such trial, the function AppUnion is called $2\ell$ times, twice for each level $1\leq i\leq\ell$. Let us now define some events of interest.
  4. (4)
    For each $1\leq i\leq\ell$ and ${\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}}\in\{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{0}},{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{1}}\}$, let us denote by $\texttt{Accurate}\texttt{AppUnion}_{i,{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}}}$ the event that the ${\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}}^{\text{th}}$ call to AppUnion corresponding to level $i$ returns a value $\textsf{sz}^{i}_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}}}$ that satisfies 1($\frac{1}{(1+\beta)^{i}}\cdot|\mathcal{L}(P^{i}_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}}})|\leq\textsf{sz}^{i}_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}}}\leq(1+\beta)^{i}\cdot|\mathcal{L}(P^{i}_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}}})|$. Here, $P^{i}$ is the argument of the call at level $i$, and $P^{i}_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}}}$ is the ${\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}}$-predecessor of $P^{i}$. For ease of notation, we also define $\texttt{Accurate}\texttt{AppUnion}=\bigcap_{\begin{subarray}{l}1\leq i\leq\ell\\ {\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}}\in\{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{0}},{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{1}}\}\end{subarray}}\texttt{Accurate}\texttt{AppUnion}_{i,{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}}}$
  5. (5)
    Let $\mathtt{Fail}_{1}$ be the event that $\mathtt{ret}_{\textnormal{{sample}}}=\bot$ is returned because $\varphi>1$ at the time of return (Line 3). Let $\mathtt{Fail}_{2}$ be the event that $\mathtt{ret}_{\textnormal{{sample}}}=\bot$ is returned at Line 3. Finally, let $\mathtt{Fail}=\mathtt{Fail}_{1}\uplus\mathtt{Fail}_{2}$.
The following presents the correctness of Algorithm 3, while its proof is presented in Section 5.
Theorem 2.
Let $\gamma_{0},\beta$ and $\{\mathsf{S}({p}^{i})\}_{1\leq i<\ell,p\in Q}$ satisfy the following:
$\begin{array}[]{c}\gamma_{0}=\frac{2}{3e}\cdot\frac{1}{\mathsf{N}({q}^{\ell})},\quad\quad\beta\leq\frac{1}{4\cdot n^{2}},\quad\quad\text{ and,}\\ \forall\;1\leq i\leq\ell,\forall p\in Q,\;\;|\mathsf{S}({p}^{i})|\geq\frac{24\cdot e}{\beta^{2}}\log\left(\frac{16mn}{\eta}\right)\end{array}$
Then for the trial corresponding to the invocation $\textnormal{{sample}}(\ell,\{{q}^{\ell}\},\lambda,\gamma_{0},\beta,\eta)$, the following statements hold true
  1. (1)
    For each $w\in\mathcal{L}({q}^{\ell})$,
    ${\mathsf{Pr}}\!\left[\mathtt{ret}_{\textnormal{{sample}}}=w\;\;\big{|}\texttt{Accurate}\texttt{AppUnion}\cap\texttt{AccurateN}_{\ell,q}\right]=\gamma_{0}$
  2. (2)
    ${\mathsf{Pr}}\!\left[\mathtt{Fail}\mid\texttt{Accurate}\texttt{AppUnion}\cap\texttt{AccurateN}_{\ell,q}\right]\leq 1-\frac{2}{3e^{2}}$
  3. (3)
    ${\mathsf{Pr}}\!\left[\texttt{Accurate}\texttt{AppUnion}\;\;\Bigg{|}\;\ \texttt{AccurateN}_{\leq\ell-1};{{\mathsf{U}}_{\leq\ell-1}}\right]\geq 1-\frac{\eta}{2}$
Input : NFA $\mathcal{A}$ with $m$ states, $n\in\mathbb{N}$ presented in unary
Parameters : Accuracy parameter $\varepsilon$ and confidence parameter $\delta$
1 $\beta\leftarrow\frac{\varepsilon}{4n^{2}}$ $\eta\leftarrow\frac{\delta}{2\cdot n\cdot m}$ $\mathtt{ns}\leftarrow\frac{4096\cdot e\cdot n^{4}}{\epsilon^{2}}\log\left(\frac{4096\cdot m^{2}\cdot n^{2}\cdot\log(\varepsilon^{-2})}{\delta}\right)$ $\mathtt{xns}\leftarrow\mathtt{ns}\cdot 12\cdot(1-\frac{2}{3e^{2}})^{-1}\cdot\log(8/\eta)$ Construct the labeled directed acyclic graph $\mathcal{A}_{\textsf{unroll}}$ from $\mathcal{A}$ input $n$ for $\ell\in\{0,\ldots,n\}\text{ and }q\in Q$ do
2     if $\ell=0$ then
3          $\big{(}\mathsf{N}({q}^{\ell}),\mathsf{S}({q}^{\ell})\big{)}\leftarrow(q\in F)$ $\big{(}1,\{\lambda\}\big{)}$ : $\big{(}0,\varnothing)$
4     foreach ${\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}}\in\{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{0}},{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{1}}\}$ do
5          $\{p_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}},1},\ldots,p_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}},k_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}}}}\}\leftarrow\textsf{Pred}(q,{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}})$ $\beta^{\prime}=(1+\beta)^{\ell-1}-1$ $\textsf{sz}_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}}}\leftarrow\texttt{AppUnion}_{\beta,\eta/2}\bigg{(}\beta^{\prime},$ $(\mathsf{S}({p_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}},1}}^{\ell-1}),\mathsf{N}({p_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}},1}}^{\ell-1})),\ldots,(\mathsf{S}({p_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}},k_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}}}}}^{\ell-1}),\mathsf{N}({p_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}},k_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}}}}}^{\ell-1})),$ $\mathcal{A}_{\textsf{unroll}}\bigg{)}$   // The last parameter $\mathcal{A}_{\textsf{unroll}}$ is a proxy for the membership oracles required by AppUnion
6
7    $\mathsf{N}({q}^{\ell})\leftarrow\textsf{sz}_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{0}}}+\textsf{sz}_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{1}}}$ $\mathsf{S}({q}^{\ell})\leftarrow\varnothing$ repeat $\mathtt{xns}$ times:
8          if $|\mathsf{S}({q}^{\ell})|<\mathtt{ns}$ then
9               $w\leftarrow$ sample(,$\ell,\{{q}^{\ell}\},\lambda,\frac{2}{3e}\cdot\frac{1}{\mathsf{N}({q}^{\ell})},\beta,\frac{\eta}{2\cdot\mathtt{xns}}$) if $w\neq\bot$ then
10                    $\mathsf{S}({q}^{\ell})\leftarrow\mathsf{S}({q}^{\ell})\boldsymbol{\!\cdot}\!\mathsf{append}(w)$
11
12          else break
13    let $\alpha_{{q}^{\ell}}=\mathtt{ns}{-}|\mathsf{S}({q}^{\ell})|$ if $\alpha_{{q}^{\ell}}>0$ then
14          let $w_{{q}^{\ell}}$ be some word in $\mathcal{L}({q}^{\ell})$ repeat $\alpha_{{q}^{\ell}}$ times:  $\mathsf{S}({q}^{\ell}){\leftarrow}\mathsf{S}({q}^{\ell})\boldsymbol{\!\cdot}\!\mathsf{append}(w_{{q}^{\ell}})$
15
return $\mathsf{N}({q}^{n}_{F})$
Algorithm 4 Algorithm for Estimating $|\mathcal{L}(\mathcal{A})_{n}|$
The main algorithm. Algorithm 4 first constructs (in Line 4) the labeled directed acyclic graph $\mathcal{A}_{\textsf{unroll}}$. The algorithm then goes over all the states and all levels $0\leq\ell\leq n$ to construct the sampled sets and the size estimates inductively. In Lines 4-4, the algorithm caters for the base case ($\ell=0$). In the inductive case (Lines 4-4) the algorithm computes the estimate $\textsf{sz}_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{1}}}$ and $\textsf{sz}_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{0}}}$, where $\textsf{sz}_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}}}$ is the estimate of the size of $\cup_{p_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}},i}\in\textsf{Pred}(q,{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}})}\mathcal{L}({p_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}},i}}^{\ell-1})$. The size estimates of, and a sample subsets of the sets $\{\mathcal{L}({q}^{\ell-1})\}_{q\in Q}$ are available inductively at this point. Further, membership in these languages can be easily checked. As a result, the algorithm uses the subroutine AppUnion to compute the size of the union (in Line 4). Once the size estimates $\textsf{sz}_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{0}}}$ and $\textsf{sz}_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{1}}}$ are determined, the algorithm constructs the sampled subset of $\mathcal{L}({q}^{\ell})$ by making $\mathsf{xns}$ calls to the subroutine sample (Lines 4-4). Theorem 3 presents the desired correctness statement of our FPRAS  Algorithm 4, and its proof is presented in Section 6.
Theorem 3.
Given an NFA $\mathcal{A}$ with $m$ states and $n\in\mathbb{N}$ (presented in unary), Algorithm 4 returns $\mathsf{Est}$ such that the following holds:
${\mathsf{Pr}}\!\left[\frac{|L(F^{n})|}{1+\epsilon}\leq\mathsf{Est}\leq(1+\epsilon)|L(F^{n})|\right]\geq 1-\delta$
Moreover, the algorithm has time complexity $\widetilde{O}(m^{4}n^{11}\cdot\frac{1}{\epsilon^{4}}\cdot\log^{2}(1/\delta))$, where the tilde hides for polynomial factors of $\log(m+n)$.
4. Correctness of Algorithm 2
The proof is in the same lines as that of Karp and Luby (Karp and Luby, 1985); for completeness, we present a full proof.
Proof.
Let $\mathsf{Fail}$ be the event that the output of the algorithm is not within $\frac{|\cup_{i=1}^{k}T_{i}|}{(1+\epsilon)(1+\epsilon_{\textsf{sz}})}$ and $|\cup_{i=1}^{k}T_{i}|{(1+\epsilon)(1+\epsilon_{\textsf{sz}})}$. Clearly, we need to upper bound $\Pr[\mathsf{Fail}]$ by $\delta$.
Note that the algorithm in each of the $t$ runs of the loop (lines 2-2) tries to draw an element from some $S_{i}$. So if we assume that the size of the sets $S_{i}$ for all $i$ is more than $t$, then the condition is the if (in line 2 will be satisfied and in that case the else clause in line 2 to 2 is redundant.
We will prove the correctness of  Algorithm 2 in two parts. In the first part we will prove that if the size of the sets $S_{i}$ is greater than $t$ then $\Pr[\mathsf{Fail}]\leq\frac{\delta}{2}$. In the second part we will prove that if the sets $S_{i}$ has size thresh (much smaller than $t$) then the probability that the algorithm will ever need the else clause in line 2 to 2 is less that $\delta/2$. The two parts combined proves that if the sets $S_{i}$s are of size threshthen $\Pr[\mathsf{Fail}]\leq\delta$ which is what is claimed in the theorem.
For all $\sigma\in\cup_{i=1}^{k}T_{i}$ let $L(\sigma)$ denote the smallest index $i$ such that $\sigma\in T_{i}$. In other words, $\sigma$ is in $T_{L(\sigma)}$ and for all $i<L(\sigma)$, $\sigma\not\in T_{i}$. Observe $\mathcal{U}_{1}:=\left\{(\sigma,L(\sigma)\ \mid\ \sigma\in\cup_{i=1}^{k}T_{i}\right\}$ has size exactly $\left|\cup_{i=1}^{k}T_{i}\right|$. Furthermore, $\mathcal{U}_{2}:=\left\{(\sigma,i)\ \mid\ \sigma\in T_{i}\right\}$ has size $\sum_{i=1}^{k}|T_{i}|$.
Part 1: If $\forall i$, $|S_{i}|\geq t$ then $\Pr[\mathsf{Fail}]\leq\frac{\delta}{2}$.
Note that in this case the condition in the if clause in line 2 is always satisfied. For a run of the loop (lines 2-2), consider the lines 2 to 2. We say the pair $(\sigma,i)$ is sampled if in one of the iteration of the loop $i$ is sampled in line 2 and then in that same iteration $\sigma$ is obtained from $S_{i}.\mathsf{deque()}$ in the line 2.
For any $i_{0}\in[t]$ and $\sigma_{0}\in S_{i_{0}}$, we have
[…]
$=$ $\frac{\textsf{sz}_{i_{0}}}{\sum_{i}^{k}\textsf{sz}_{i}}\times\frac{1}{|T_{i_{0}}|}$
The last equality is because we have assumed $\forall i$, $|S_{i}|\geq t$, so the if condition in line 2 is always satisfied. Hence once $i_{0}$ is sampled in line 2 probability that $\sigma_{0}$ is obtained is exactly $1/|T_{i_{0}}|$, since the set $S_{i_{0}}$ contains elements sampled uniformly (with replacement) from $T_{i_{0}}$. Therefore,
$\frac{1}{(1+\epsilon_{\textsf{sz}})\sum_{i=1}^{k}\textsf{sz}_{i}}\leq\Pr[(\sigma_{0},i_{0})\mbox{ is sampled }]\leq(1+\epsilon_{\textsf{sz}})\frac{1}{\sum_{i=1}^{k}\textsf{sz}_{i}}.$
Let $Y_{j}$ be the random variable denoting whether the counter $Y$ is increased in the $j$th iteration of the loop. Note that $Y_{j}=1$ if the pair $(\sigma_{0},i_{0})$ sampled in the $j$th iteration is in $\mathcal{U}_{1}$. This is because in line 2 it is checked if the sampled pair is in $\mathcal{U}_{1}$.
So, $\mathbb{E}[Y_{j}]=\sum_{(\sigma,i)\in\mathcal{U}_{1}}\Pr[(\sigma,i)\mbox{ is sampled}]$. Thus,
$\frac{|\mathcal{U}_{1}|}{(1+\epsilon_{\textsf{sz}})\sum_{i=1}^{k}\textsf{sz}_{i}}\leq\mathbb{E}[Y_{j}]\leq(1+\epsilon_{\textsf{sz}})\frac{|\mathcal{U}_{1}|}{\sum_{i=1}^{k}\textsf{sz}_{i}}$
Note that $Y=\sum_{i}^{t}Y_{i}$. Thus at the end of the algorithm the expected value of $Y$ is between $\frac{t|\mathcal{U}_{1}|}{(1+\epsilon_{\textsf{sz}})\sum_{i=1}^{k}\textsf{sz}_{i}}$ and $(1+\epsilon_{\textsf{sz}})^{2}\frac{t|\mathcal{U}_{1}|}{\sum_{i=1}^{k}\textsf{sz}_{i}}$.
Now by Chernoff bound we have
$\Pr[Y\not\in(1\pm\frac{\epsilon}{2})\mathbb{E}[Y]]\leq 2exp(-\frac{\epsilon^{2}\mathbb{E}[Y]}{12}).$
We also know that,
$\mathbb{E}[Y]\geq$ $\frac{t|\mathcal{U}_{1}|}{(1+\epsilon_{\textsf{sz}})\sum_{i=1}^{k}\textsf{sz}_{i}}\geq\frac{t\max_{i}|T_{i}|}{(1+\epsilon_{\textsf{sz}})\sum_{i=1}^{k}\textsf{sz}_{i}}\geq\frac{t\max_{i}\textsf{sz}_{i}}{(1+\epsilon_{\textsf{sz}})^{2}\sum_{i=1}^{k}\textsf{sz}_{i}}$
$\geq$ $12\log(\frac{4}{\delta})$
The last inequality is due to the setting of the parameter $t$. Thus, by Chernoff bound, probability that the output of the algorithm is not between $(1-\frac{\epsilon}{2})\frac{|\mathcal{U}_{1}|}{(1+\epsilon_{\textsf{sz}})}$ and $(1+\frac{\epsilon}{2})(1+\epsilon_{\textsf{sz}})|\mathcal{U}_{1}|$ is at most $\delta/2$. Since $(1-\frac{\epsilon}{2})\geq\frac{1}{1+\epsilon}$, so we have $\Pr[\mathsf{Fail}\ \cap\ \overline{\mathsf{Bad}}]\leq\frac{\delta}{2}$.
Part 2: If for all $i$, $|S_{i}|\geq\texttt{thresh}$
Note that, the else clause in line 2 to 2 is ever reached, if for some $i$, the number of times $i$ is sampled in the $t$ iterations is more than thresh.
Let $\mathsf{Bad}_{i}$ denote the event that the index $i$ is sampled (in line 2) more that thresh number of times during the course of the algorithm, that is during the $t$ runs of the loop (lines 2-2). And let the event $\mathsf{Bad}$ be $\cup_{i=1}^{k}\mathsf{Bad}_{i}$. We will now upper bound, $\Pr[\mathsf{Bad}_{i}]$. Let $W_{i}$ be the random variable counting the number times $i$ is sampled. $\mathbb{E}[W_{i}]=t\times\frac{\textsf{sz}_{i}}{\sum_{i=1}^{k}\textsf{sz}_{i}}$ which is less than $\frac{12(1+\epsilon_{\textsf{sz}})^{2}}{\epsilon^{2}}\log(4/\delta)$. By simple application of Chernoff Bound we see that
$\Pr[W_{i}\geq(1+\Delta)\mathbb{E}[W_{i}]]\leq exp(-\frac{\Delta^{2}\mathbb{E}[W_{i}]}{2+\Delta})\leq exp(-\frac{\Delta\mathbb{E}[W_{i}]}{2}),$
the last inequality is for $\Delta\geq 1$. For our purpose if we put $\Delta=\log(\frac{2k}{\delta})\frac{1}{\mathbb{E}[W_{i}]}$, since $|S_{i}|\geq\texttt{thresh}\geq\frac{12\cdot(1+\epsilon_{\textsf{sz}})^{2}}{\varepsilon^{2}}\cdot\log(\frac{4}{\delta})$ + $\log(\frac{2k}{\delta})$, we have $\Pr[\mathsf{Bad_{i}}]\leq\frac{\delta}{2k}$. So by union bound $\Pr[\mathsf{Bad}]\leq\delta/2$.
5. Correctness of Algorithm 3
In the following three sections, we prove the three parts of Theorem 2 individually.
5.1. Proof of Theorem 2(1)
Consider an execution of Algorithm 3 and let $w=w_{1}w_{2}\ldots w_{\ell}\in\mathcal{L}({q}^{\ell})$ be the string constructed right before entering the branch at Line 3. Let $p_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{0}},i}$ be the value of the variable $\mathsf{pr_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{0}}}}$ at level $i$ and let $v$ be the value of $\varphi$ before the function returns. Then we have,
$v=\frac{\gamma_{0}}{p_{w_{\ell},\ell}\cdot p_{w_{\ell-1},\ell-1}\cdots p_{w_{1},1}}$
Let $\textsf{sz}^{\ell}_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{0}}},\textsf{sz}^{\ell}_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{1}}},\textsf{sz}^{\ell-1}_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{0}}},\textsf{sz}^{\ell-1}_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{1}}},\ldots\textsf{sz}^{1}_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{0}}},\textsf{sz}^{1}_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{1}}}$ be the estimates obtained in the $2\ell$ calls to AppUnion during the run of $\textnormal{{sample}}()$. Then,
$p_{w_{i},i}=\frac{\textsf{sz}^{i}_{w_{i}}}{\textsf{sz}^{i}_{w_{i}}+\textsf{sz}^{i}_{1-w_{i}}}$
Thus,
(2) $v=\frac{\gamma_{0}}{\prod\limits_{i=1}^{\ell}\frac{\textsf{sz}^{i}_{w_{i}}}{\textsf{sz}^{i}_{w_{i}}+\textsf{sz}^{i}_{1-w_{i}}}}=\gamma_{0}\cdot\prod\limits_{i=1}^{\ell}\frac{\textsf{sz}^{i}_{w_{i}}+\textsf{sz}^{i}_{1-w_{i}}}{\textsf{sz}^{i}_{w_{i}}}$
Recall that $P^{i}$ is the argument of the call at level $i$, and $P^{i}_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}}}$ is the ${\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}}$-predecessor of $P^{i}$. Observe that $\mathcal{L}(P^{i})=\mathcal{L}(P^{i}_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{0}}})\uplus\mathcal{L}(P^{i}_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{1}}})$, and thus,
$|\mathcal{L}(P^{i}_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{0}}})|+|\mathcal{L}(P^{i}_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{1}}})|=|\mathcal{L}(P^{i})|$
Hence, under the event $\texttt{Accurate}\texttt{AppUnion}_{i,{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{0}}}\cap\texttt{Accurate}\texttt{AppUnion}_{i,{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{1}}}$ we have
(3) $\frac{\textsf{sz}^{i}_{w_{i}}+\textsf{sz}^{i}_{1-w_{i}}}{\textsf{sz}^{i}_{w_{i}}}\leq(1+\beta)^{2i}\cdot\frac{|\mathcal{L}(P^{i})|}{|\mathcal{L}(P^{i-1})|}$
This gives us,
$\begin{array}[]{c}v\leq\gamma_{0}\cdot\prod\limits_{i=1}^{\ell}(1+\beta)^{2i}\cdot\frac{|\mathcal{L}(P^{i})|}{|\mathcal{L}(P^{i-1})|}=\gamma_{0}{\cdot}\Bigg{(}\prod\limits_{i=1}^{\ell}(1+\beta)^{2i}\Bigg{)}{\cdot}\Bigg{(}\prod\limits_{i=1}^{\ell}\frac{|\mathcal{L}(P^{i})|}{|\mathcal{L}(P^{i-1})|}\Bigg{)}\end{array}$
Now, observe that $|\mathcal{L}(P^{0})|=1$ and $P^{\ell}=\{{q}^{\ell}\}$. Thus, under the event $\big{(}\bigcap_{\begin{subarray}{l}1\leq i\leq\ell\\ {\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}}\in\{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{0}},{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{1}}\}\end{subarray}}\texttt{Accurate}\texttt{AppUnion}_{i,{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}}}\big{)}=\texttt{Accurate}\texttt{AppUnion}$, we have
(5) $\begin{array}[]{c}v\leq\gamma_{0}{\cdot}\Bigg{(}\prod\limits_{i=1}^{\ell}(1+\beta)^{2i}\Bigg{)}{\cdot}|\mathcal{L}({q}^{\ell})|=\gamma_{0}{\cdot}(1+\beta)^{\ell(\ell+1)}{\cdot}|\mathcal{L}({q}^{\ell})|\end{array}$
Assuming the event $\texttt{AccurateN}_{\ell,q}$ also holds, we have $\frac{|\mathcal{L}({q}^{\ell})|}{\mathsf{N}({q}^{\ell})}\leq(1+\beta)^{\ell}$. Under this assumption, substituting $\gamma_{0}=\frac{2}{3e}\cdot\frac{1}{\mathsf{N}({q}^{\ell})}$, $\beta\leq\frac{1}{4\cdot n^{2}}$, and using the inequality $(1+1/4n^{2})^{n}\leq e$, we have,
$\begin{array}[]{c}v\leq\frac{2}{3e}{\cdot}(1+\beta)^{\ell(\ell+2)}\leq\frac{2}{3e}\cdot e\leq 1\\ \end{array}$
It then follows that
${\mathsf{Pr}}\!\left[\mathtt{Fail}_{1}\mid\texttt{Accurate}\texttt{AppUnion}\cap\texttt{AccurateN}_{\ell,q}\right]=0$
Substituting, $v=\frac{\gamma_{0}}{p_{w_{\ell},\ell}\cdot p_{w_{\ell-1},\ell-1}\cdots p_{w_{1},1}}$, we then have
$\begin{array}[]{l}{\mathsf{Pr}}\!\left[\mathtt{ret}_{\textnormal{{sample}}}=w\mid\texttt{Accurate}\texttt{AppUnion}\cap\texttt{AccurateN}_{\ell,q}\right]\\ =v\cdot\prod_{i=\ell}^{1}p_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{0}},i}=\frac{\gamma_{0}}{p_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{0}},\ell}\cdot p_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{0}},\ell-1}\cdots p_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{0}},1}}\cdot\prod_{i=\ell}^{1}p_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{0}},i}=\gamma_{0}\end{array}$
5.2. Proof of Theorem 2(2)
We will now estimate the (conditional) probability of the event $\mathtt{Fail}$. First, observe that for each string $w^{\prime}\notin\mathcal{L}({q}^{\ell})$, we have ${\mathsf{Pr}}\!\left[\mathtt{ret}_{\textnormal{{sample}}}=w^{\prime}\right]=0$ because any string $w$ that the algorithm outputs is such that there is a path from the initial state to ${q}^{\ell}$ labeled with $w$, and (b) the unrolled automata is such that any path from start state to some state ${p}^{i}$ corresponds to a string in $\mathcal{L}({p}^{i})$.
We observe that $\overline{\mathtt{Fail}}$ is the event that $\mathtt{ret}_{\textnormal{{sample}}}$ is a string in $\mathcal{L}({q}^{\ell})$, each of which is a disjoint event. Thus,
${\mathsf{Pr}}\!\left[\mathtt{Fail}\mid\texttt{Accurate}\texttt{AppUnion}\cap\texttt{AccurateN}_{\ell,q}\right]$
$=1{-}\sum\limits_{w\in\mathcal{L}({q}^{\ell})}{\mathsf{Pr}}\!\left[\mathtt{ret}_{\textnormal{{sample}}}=w\mid\texttt{Accurate}\texttt{AppUnion}\cap\texttt{AccurateN}_{\ell,q}\right]$
$=1{-}\sum\limits_{w\in\mathcal{L}({q}^{\ell})}\gamma_{0}=1{-}\gamma_{0}|\mathcal{L}({q}^{\ell})|\leq 1-\frac{2}{3e}\cdot\frac{|\mathcal{L}({q}^{\ell})|}{\mathsf{N}({q}^{\ell})}$
$\leq 1-\frac{2}{3e}\cdot\frac{1}{(1+\beta)^{\ell}}\leq 1-\frac{2}{3e^{2}}$
5.3. Proof of Theorem 2(3)
We would like to use Theorem 1 to prove Theorem 2(3). Towards this, we first establish that the pre-conditions of Theorem 1 hold.
  1. (1)
    For $i\leq\ell$, if we substitute $\beta^{\prime}=(1+\beta)^{i-1}-1,\eta^{\prime}=\eta/{4n}$, we get $\texttt{thresh}=24\cdot\frac{(1+\beta)^{2i-2}}{\beta^{2}}\cdot\log\left(\frac{16k_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}}}n}{\eta}\right)$, where $k_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}}}$ is the number of sets in the argument of AppUnion. Observe that, $k_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}}}\leq m$ and that, for each $i\leq\ell$, we have $(1+\beta)^{2i-2}\leq(1+\beta)^{4n^{2}}\leq e$. Thus for each ${q}^{i}$, we have
    $\texttt{thresh}\leq\frac{24\cdot e}{\beta^{2}}\log\left(\frac{16mn}{\eta}\right)\leq|\mathsf{S}({q}^{i-1})|$
  2. (2)
    Since we are considering the distribution obtained when $\mathsf{S}({q}^{i})$ is replaced with $\mathsf{U}({q}^{i})$ for each $i\leq\ell-1$ (observe the ${\mathsf{Pr}}\!\left[;{{\mathsf{U}}_{\leq\ell-1}}\right]$ in the statement of Theorem 2(3)), we directly have that each input $\{S_{i}\}_{i}$ to every call to AppUnion has at least thresh elements (as established) elements obtained by repeatedly sampling $\{T_{i}\}_{i}$ uniformly.
  3. (3)
    The event $\texttt{AccurateN}_{\leq\ell-1}$ demands that for each $i\leq\ell-1$, we have $\frac{|\mathcal{L}({q}^{i})|}{(1+\beta)^{i}}\leq\mathsf{N}({q}^{i})\leq(1+\beta)^{i}|\mathcal{L}({q}^{i})|$.
From Theorem 1, it thus follows that, for all $i$,
${\mathsf{Pr}}\!\left[\texttt{Accurate}\texttt{AppUnion}_{i,{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}}}\;\;\Bigg{|}\;\ \texttt{AccurateN}_{\leq\ell-1};{{\mathsf{U}}_{\leq\ell-1}}\right]\geq 1-\eta^{\prime}$
Therefore, applying union-bound, we have
$\begin{array}[]{l}{\mathsf{Pr}}\!\left[\bigcap_{\begin{subarray}{l}1\leq i\leq\ell\\ {\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}}\in\{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{0}},{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{1}}\}\end{subarray}}\texttt{Accurate}\texttt{AppUnion}_{i,{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}}}\;\;\Bigg{|}\;\ \texttt{AccurateN}_{\leq\ell-1};{{\mathsf{U}}_{\leq\ell-1}}\right]\\ \geq 1-2n\cdot\eta^{\prime}=1-\frac{\eta}{2}\end{array}$
6. Correctness of Algorithm 4
We first establish some preliminaries relevant to our proof.
6.1. Total Variation Distance
The total variation distance is a statistical distance between two discrete valued random variables $X,Y$ over domain $\Omega$ and is defined as follows:
$\mathsf{TV}(X,Y)$ $=$ $\frac{1}{2}\sum_{v\in\Omega}\left|\Pr[X=v]-\Pr[Y=v]\right|$
$=$ $\sum_{v\,:\,\Pr[X=v]\geq\Pr[Y=v]}\left(\Pr[X=v]-\Pr[Y=v]\right)\;.$
We will use the following folklore result (c.f. Lemma 6.3 of  (Vadhan, 2012)). Here, for a random variable $X$ over the sample space $\Omega$, we use the notation ${\mathsf{Pr}}\!\left[E:X\right]$ to denote the probability of the event $E\subseteq\Omega$.
Lemma 6.0.
Let $X$ and $Y$ be the random variables taking values over a measurable set $\Omega$. Let $f:\Omega\to\hat{\Omega}$ be a (possibly randomized) function with measurable co-domain $\hat{\Omega}$. Let $E\subseteq\hat{\Omega}$ be an arbitrary event. Then, the following holds:
$|{\mathsf{Pr}}\!\left[E:{f(X)}\right]-{\mathsf{Pr}}\!\left[E:f(Y)\right]|\leq\mathsf{TV}(X,Y)$
6.2. Proof of Theorem 3
We prove the main result (Theorem 3) by induction on the level $\ell$. For each level $\ell$, we first characterize the accuracy of the computed estimates $\{\mathsf{N}({q}^{\ell})\}_{q\in Q}$ and the quality of the sampled sets $\{\mathsf{S}({q}^{\ell})\}_{q\in Q}$, assuming that these were computed when the samples at the lower levels $11,2,\ldots,\ell-$ 1 are perfectly uniform and have size thresh. After this, we characterize the real computed estimates and samples by arguing that the distance of the corresponding random variables, from those computed using perfectly uniform samples, is small. We will be using the total variation distance and Lemma 1 to structure the proof. We first establish some helpful auxiliary results (Lemma 2 and Lemma 3) towards our main proof. As before, we will often invoke the the probability space obtained by replacing the sets $\mathsf{S}({q}^{j})$ by $\mathsf{U}({q}^{j})$ for all $q\in Q$ and for all $1\leq j\leq i$. In addition to the previously introduced notation ${\mathsf{Pr}}\!\left[;{{\mathsf{U}}_{\leq i}}\right]$, we will also use the notation ${\mathsf{TV}}(X,Y\;;{\mathsf{U}}_{\leq i})$ (resp. ${\mathsf{TV}}(X,Y\;\;\big{|}\;\;\mathcal{F}\;;{\mathsf{U}}_{\leq i})$) to denote the total variational distance between the distribution induced by random variables $X,Y$ (resp. distribution induced by $X,Y$ conditioned on the event $\mathcal{F}$) when the $\mathsf{S}({q}^{j})$ in Algorithm 3 and Algorithm 4 is replaced by a set $\mathsf{U}({q}^{j})$ obtained by unirofmly sampling (with replacement) $|\mathsf{S}({q}^{j})|$ many elements from $\mathcal{L}({q}^{j})$, for all $q\in Q$ and for all $1\leq j\leq i$.
Lemma 6.0.
For each $\ell\leq n$ and $q\in Q$, we have
${\mathsf{Pr}}\!\left[\texttt{AccurateN}_{q,\ell}\;\;\big{|}\;\;\texttt{AccurateN}_{\leq\ell-1};{{\mathsf{U}}_{\leq\ell-1}}\right]\geq 1-\eta$
Proof.
[…]
thresh $=24\cdot\frac{(1+\beta)^{2\ell-2}}{\beta^{2}}\cdot\log\left(\frac{4k_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}}}}{\eta^{\prime}}\right)\leq\frac{24\cdot e}{\beta^{2}}\log\left(\frac{8m^{2}n}{\delta}\right)$
$\leq\frac{384\cdot e\cdot n^{4}}{\epsilon^{2}}\log\left(\frac{8m^{2}}{\delta}\right)\leq\mathtt{ns}$
Therefore, from Theorem 1, we have
${\mathsf{Pr}}\!\left[\begin{array}[]{l}\frac{|L(\textsf{Pred}(q^{\ell},{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}}))|}{(1+\beta)^{\ell}}\leq\textsf{sz}_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}}}\\ \quad\quad\leq(1+\beta)^{\ell}\;|L(\textsf{Pred}(q^{\ell},{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}}))|\end{array}\big{|}\;\;\texttt{AccurateN}_{\ell-1};{{\mathsf{U}}_{\ell-1}}\right]\geq 1-\frac{\eta}{2}$
Since $\mathsf{N}({q}^{\ell})=\textsf{sz}_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{0}}}+\textsf{sz}_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{1}}}$, we have
${\mathsf{Pr}}\!\left[\begin{array}[]{l}\frac{|\mathcal{L}({q}^{\ell})|}{(1+\beta)^{\ell}}\leq\mathsf{N}({q}^{\ell})\\ \quad\quad\leq(1+\beta)^{\ell}|\mathcal{L}({q}^{\ell})|\end{array}\mid\;\;\texttt{AccurateN}_{\leq\ell-1};{{\mathsf{U}}_{\leq\ell-1}}\right]\geq 1-\eta$
Lemma 6.0.
For each $\ell\leq n$ and $q\in Q$, we have
${\mathsf{TV}}(S(q^{\ell}),U(q^{\ell})\;\;\big{|}\;\;\texttt{AccurateN}_{\leq\ell}\;;{\mathsf{U}}_{\ell-1})\leq\eta$
Proof.
Fix $\ell$ and $q$. Let $\mathtt{Fail}^{j}$ denote the event that the $j^{\text{th}}$ call to sample (in the loop starting at Line 4) for $i=j$ returns $\bot$. Similarly, let $\texttt{Accurate}\texttt{AppUnion}^{j}$ denote the event AccurateAppUnion corresponding to the call to sample for $i=j$. Using Iversion bracket notation, we denote the indicator variable $[\mathtt{Fail}^{j}]$. Also, we use SmallS to denote the event that $\alpha_{{q}^{\ell}}>0$. Observe that SmallS occurs if and only if $\sum\limits_{j=1}^{\mathtt{xns}}[\mathtt{Fail}^{j}]\geq\mathtt{xns}-\mathtt{ns}$.
Recall, we have, for each $j$
${\mathsf{Pr}}\!\left[[\mathtt{Fail}^{j}]\;\;\big{|}\;\;\texttt{Accurate}\texttt{AppUnion}^{j}\cap\texttt{AccurateN}_{\leq\ell};{{\mathsf{U}}_{\leq\ell-1}}\right]\leq 1-\frac{2}{3e^{2}}$
Therefore, from Chernoff bound, we have
${\mathsf{Pr}}\!\left[\textsc{SmallS}\;\;\big{|}\;\;\bigcap\limits_{j}\texttt{Accurate}\texttt{AppUnion}^{j}\cap\texttt{AccurateN}_{\leq\ell};{{\mathsf{U}}_{\leq\ell-1}}\right]\leq\frac{\eta}{2}$
Furthermore, from Theorem 2 (part 3), we have
${\mathsf{Pr}}\!\left[\overline{\bigcap\limits_{j}\texttt{Accurate}\texttt{AppUnion}^{j}}\big{|}\;\;\texttt{AccurateN}_{\leq\ell};{{\mathsf{U}}_{\leq\ell-1}}\right]\leq\eta/2$
Note that assuming $\texttt{AccurateN}_{\leq\ell}$ and ${\mathsf{U}}_{\leq\ell-1}$, is the event SmallS does not happen and for all $j$ $\texttt{Accurate}\texttt{AppUnion}^{j}$ holds then the distribution of $S(q^{\ell})$ and $U(q^{\ell})$ is identical. So,
$\mathsf{TV}(S(q^{\ell}),U(q^{\ell})\;\;\big{|}\;\;\texttt{AccurateN}_{\leq\ell}\;;\;{\mathsf{U}}_{\leq\ell-1})$
$\leq{\mathsf{Pr}}\!\left[\textsc{SmallS}\cup\overline{\bigcap\limits_{j}\texttt{Accurate}\texttt{AppUnion}^{j}}\;\;\big{|}\;\;\texttt{AccurateN}_{\leq\ell};{{\mathsf{U}}_{\leq\ell-1}}\right]$
$\leq{\mathsf{Pr}}\!\left[\textsc{SmallS}\;\;\big{|}\bigcap\limits_{j}\texttt{Accurate}\texttt{AppUnion}^{j}\cap\texttt{AccurateN}_{\leq\ell};{{\mathsf{U}}_{\leq\ell-1}}\right]$
$\quad+{\mathsf{Pr}}\!\left[\overline{\bigcap\limits_{j}\texttt{Accurate}\texttt{AppUnion}^{j}}\;\;\big{|}\;\;\texttt{AccurateN}_{\leq\ell};{{\mathsf{U}}_{\leq\ell-1}}\right]$
$\leq\eta/2+\eta/2=\eta$
Now, we are ready to state and prove the main theorem.
Proof of Theorem 3.
For ease of exposition, we will refer to the event $\texttt{AccurateN}_{n}$ in our analysis. Applying  Lemma 1, we have
${\mathsf{Pr}}\!\left[\overline{\texttt{AccurateN}_{n}}\right]\leq$
$\sum_{\ell\in[1,n];q\in Q}{\mathsf{Pr}}\!\left[\overline{\texttt{AccurateN}_{q,\ell}}\;\;\big{|}\;\;\texttt{AccurateN}_{\leq\ell-1};{{\mathsf{U}}_{\leq\ell-1}}\right]$
$\qquad+\sum\limits_{\ell\in[1,n-1];q\in Q}\mathsf{TV}(S(q^{\ell}),U(q^{\ell})\mid\texttt{AccurateN}_{\leq\ell}\;;{\mathsf{U}}_{\leq\ell-1})$
By Lemma 2 and Lemma 3 and applying union bound oven all $\ell\in[1,n]$ and all $q\in Q$ we have
${\mathsf{Pr}}\!\left[\overline{\texttt{AccurateN}_{n}}\right]\leq n\cdot m\cdot\eta+n\cdot m\cdot\eta\leq\delta.$
To calculate the time complexity of Algorithm 4 we first observe that the for loop goes over $nm$ elements. Now for each $\ell$ and $q$ there is one call to $\texttt{AppUnion}_{\beta,\eta}$ with the number of sets being $O(m)$ and $\mathtt{xns}$ calls to sample(). So by Theorem 1 the time taken by AppUnion is $O(m/\beta^{2}\log(m/\eta))$ number of calls to the oracle. So the total complexity is $\widetilde{O}(mn^{4}/\epsilon^{2}\log(1/\delta))$, ignoring the factors of $\log(n+m)$. The sample() on the other hand is a recursive function which calls $\texttt{AppUnion}_{\beta,\eta}()$ every recursion and the recursion happens $O(n)$ times. So the time complexity for each of the $\mathtt{xns}$ calls to sample() is $\widetilde{O}(mn^{5}/\epsilon^{2}\log(1/\delta))$ calls to the oracle. Now note that the membership oracle takes $O(m^{2}n)$ steps for each call. So in total the time complexity of the Algorithm 4 is $\widetilde{O}(mn\cdot\mathtt{xns}\cdot\frac{mn^{5}}{\epsilon^{2}}\log(1/\delta)\cdot(m^{2}n))=\widetilde{O}(m^{4}n^{11}\cdot\frac{1}{\epsilon^{4}}\cdot\log^{2}(1/\delta))$.
7. Conclusions and Future Work
We consider the approximate counting problem #NFA and propose a fully polynomial time randomized approximation scheme (FPRAS) that significantly improves the prior FPRAS (Arenas et al., 2019). Our algorithm deeply exploits recent progress in the streaming algorithms literature. Given the wide range of applications of counting and sampling from the language of an NFA in program analysis, testing and more broadly in Computer Science, we envision that further improvements in the complexity of approximating #NFA is a worthwhile avenue for future work.
Acknowledgements.
The authors thank Pierre Snellart and Antoine Amarilli for introducing us to the #NFA problem, and are grateful to N.V. Vinodchandran and Pavan Aduri for insightful discussions on the Karp-Luby method for the union of sets. This work was supported in part by National Research Foundation Singapore under its NRF Fellowship Programme[NRF-NRFFAI1-2019-0004 ], Ministry of Education Singapore Tier 2 grant MOE-T2EP20121-0011, and Ministry of Education Singapore Tier 1 grants [R-252-000-B59-114] and [R-252-000-C28-133].